OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (amistre64):

well, what are the first 2 terms?

OpenStudy (amistre64):

f(1) = 0 f(2) = 1 let n=3 f(3) = f(3-1) + f(3-2) = f(2) + f(1) = 1 + 0 let n=4 and work the same process, then let n=5 and do the exact same process.

OpenStudy (anonymous):

wut do u think tha answer is @jboy_52

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

wut grade is this

OpenStudy (anonymous):

9th lol